a) How many number of rules can be made from the dataset which is given in Table 3?
b) Identify those item sets from which rules can be made. Also define the criteria of selection
c) Calculate the Support value of {Ketchup and Soda} and find the confidence value of all rules which can be possible by this item set.

Table 3.
Transaction ID Items
1 Ketchup, Hamburgers, Soda
2 Cereal, Milk, Diapers, Bread
3 Hot dogs, ketchup, Soda, Milk
4 Greeting card, Cake, Soda
5 Greeting Card, Cake, Milk, Cereal
1
Expert's answer
2021-09-06T13:17:30-0400

The transaction table is given below :

Transaction ID Items:

1 Ketchup, Hamburgers, Soda

2 Cereal, Milk, Diapers, Bread

3 Hot dogs, ketchup, Soda, Milk

4 Greeting card, Cake, Soda

5 Greeting Card, Cake, Milk, Cereal

-------------------------------------------------------------------------

The frequency of occurrence of items -

Ketchup - 2

Hamburgers - 1

Soda - 3

Cereal -2

Milk - 3

Diapers - 1

Bread - 1

Hot dogs -1

Greeting Card - 2

Cake - 2

----------------------------------------------------------------------------

S= (Soda,Milk) /T = 3/5 = 0.6

C= (Ketchup,Cereal,Greeting card,Cake)/T = 2/6 = 0.33

l= (Hamburgers,Diapers, Bread, Hot dogs )/ T= 1/6 = 0.166